What is the pH of a solution that contains 25 grams of sodium hydroxide(NaOH) dissolved in 1.2 liters of water?
1
Expert's answer
2020-06-03T15:16:47-0400

Cm=n/V= 25/1.2 x 40 =0.521 M.

pOH= -lg[0.521]= 0.283.

pH=14-pOH= 14-0.283 = 13.717.